If a rectangular computer screen is 17 inches wide and 12 inches high, then what is its diagonal measure to the nearest inch?
Answer by addingup(3677)

17^2+12^2 = h^2; sqrt h^2 = diagonal (hypotenuse)
289+144 = 433; sqrt433 = hypotenuse - Use your calculator, find the square root of 433 to get your answer